- A 5C charge is positioned at x = 0cm and a -2C charge at x = 3cm. Calculate the magnitude of the force felt by each charge and determine if the force is repulsive or attractive. Given that Coulombs constant k = 9x109 Nm2C -2


1
Expert's answer
2021-08-07T03:32:35-0400

The charges have opposite signs, so the force will be attractive. According to the third Newton's law, the moduli of forces are equal, but the directions of forces will be opposite.

The modulus of force will be

"F = k\cdot \dfrac{q_1q_2}{r^2}," where r is the distance between the charges. r is "3 - 0 = 3\,\mathrm{cm} = 0.03\,\mathrm{m}," so the force will be

"F = 9\cdot10^9\,\mathrm{Nm^2 C^{-2}} \cdot \dfrac{5\cdot(-2)}{(0.03\,\mathrm{m})^2} = 10^{14}\,\mathrm{N}."